# Mid-Infrared Dust Evolution and Late-time Circumstellar Medium Interaction in SN 2017eaw

### v1

@@ 1 Introduction
Observations of high redshift galaxies have revealed significant amounts of dust in the early universe (D. P. Marrone et al., 2018; T. Hashimoto et al., 2019; J. Witstok et al., 2023; V. Markov et al., 2024; A. Nanni et al., 2025). The majority of this dust is likely associated with core-collapse supernovae (CCSNe; H. L. Morgan & M. G. Edmunds, 2003; R. Maiolino et al., 2004; C. Gall et al., 2011; R. Schneider & R. Maiolino, 2024). Models of high redshift supernovae (SNe) and star formation rates indicate that SNe would need to produce between M⊙ of dust per SN (P. Todini & A. Ferrara, 2001; A. Sarangi et al., 2018; R. Schneider & R. Maiolino, 2024) and observations of nearby SN remnants have revealed dust masses within this range (L. Dunne et al., 2003; H. L. Morgan et al., 2003; J. Rho et al., 2008; I. De Looze et al., 2017; J. Chastenet et al., 2022; F. D. Priestley et al., 2022). However, the vast majority of near- and mid-infrared studies of nearby SNe undertaken prior to JWST have revealed significantly lower dust masses (10-2 ; C. Gall et al., 2011; T. Szalai & J. Vinkó, 2013; T. Szalai et al., 2019a; S. Tinyanont et al., 2016). These previous studies therefore suggest that infrared (IR) observations of CCSNe in the decades after explosion may be missing a significant portion of the dust.
Dust formation in SNe likely occurs in the expanding ejecta interior to the reverse shock and/or in a cold dense shell between the forward and reverse shock created by the interaction between the forward shock and surrounding dense circumstellar material (CSM) (e.g. M. Pozzo et al., 2004; S. Mattila et al., 2008; N. Smith et al., 2008, 2009). In the decades to centuries following the SN, some of the dust in the interior ejecta will be destroyed by interaction with the reverse shock; however, some percentage is expected to survive this interaction and facilitate further dust formation. There are some indications that the majority of newly formed dust in the first years post-explosion in non-interacting type II supernovae (SNe II111We use SNe II to refer only to SNe IIP/L SNe and exclude SNe IIb and IIn. However, we do include peculiar SNe II, such as SN 1987A-like objects, in this class.) is interior and optically thick and thus only visible in the spectral line profiles and not in infrared images. Studies modeling the optical nebular spectra of SN II recover higher dust masses than indicated by infrared photometry alone (M. Niculescu-Duvaz et al., 2022; S. Zsíros et al., 2024). Further, the significant dust mass in SN 1987A is only observable in the far-IR and sub-millimeter (P. Bouchet et al., 2006; M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015; P. Cigan et al., 2019). Thus even in the case where the dust is optically thin, any newly-formed dust may be too cold to detect in the near/mid-infrared.
@@
The launch of JWST has ushered in a new era of SN dust studies. JWST’s sensitivity, resolution and wavelength coverage has allowed for observations that probe both hot and cool dust around the SNe in the years following collapse. In concert with multi-wavelength ground- and space-based observations, numerous studies have utilized JWST data to constrain the dust formation around some of the nearest SNe (R. G. Arendt et al., 2023; G. Hosseinzadeh et al., 2023b; O. C. Jones et al., 2023; M. Shahbandeh et al., 2023; P. Bouchet et al., 2024; S. Gomez et al., 2024; M. Matsuura et al., 2024; M. Shahbandeh et al., 2024a, b; S. Zsíros et al., 2024; G. C. Clayton et al., 2025; A. Sarangi et al., 2025; T. Szalai et al., 2025; S. Tinyanont et al., 2025). This growing dataset is vital to understanding the onset of dust production in CCSNe and its impact on the evolution of the early Universe.
@@ 2 Observations
SN 2017eaw was discovered on 2017 May 14 in NGC 6946, a nearby galaxy ( Mpc according to the latest TRBG distance222https://edd.ifa.hawaii.edu/get_cmd.php?pgc=65001, this distance is also used in M. Shahbandeh et al. (2023)) with a high SN rate (see Figure 1). SN 2017eaw has extensive multi-wavelength observations, both pre- and post-explosion (explosion on 57885.2 MJD S. D. Van Dyk et al., 2019), and has been the subject of numerous studies (e.g. D. Y. Tsvetkov et al., 2018; J. Rho et al., 2018; C. D. Kilpatrick & R. J. Foley, 2018; S. Tinyanont et al., 2019; S. D. Van Dyk et al., 2019; T. Szalai et al., 2019b; L. Rui et al., 2019; R. J. Buta & W. C. Keel, 2019; K. E. Weil et al., 2020; M. Shahbandeh et al., 2023; K. A. Bostroem et al., 2023).
There is significant evidence of ejecta-CSM interaction in SN 2017eaw. Pre-explosion Spitzer Space Telescope (hereafter Spitzer) images show the progenitor star was surrounded by a dusty shell at 4000 (C. D. Kilpatrick & R. J. Foley, 2018). Early time detections of SN 2017eaw in the X-ray, UV, and radio indicate moderate interaction between the SN shock and the CSM (T. Szalai et al., 2019b). Further, the optical light curve of SN 2017eaw exhibits a bump peaking around a week post-explosion that further suggests early time CSM interaction (T. Szalai et al., 2019b). Years after explosion there remain signs of ongoing ejecta-CSM interaction. Hubble Space Telescope (HST) near-UV imaging reveals the SN is still UV bright (as shown in Figure 1; S. D. Van Dyk et al., 2023) and late-time ( days post-explosion) optical spectra exhibit boxy line profiles indicating that the ejecta is continuing to collide with the surrounding material (K. E. Weil et al., 2020; M. Shahbandeh et al., 2023). Given the continued detection of CSM interaction, it is likely that some pre-existing dust surrounds SN 2017eaw.
@@ 2.1 JWST/MIRI
We attempt aperture photometry on the Cycle 1 and 2 images using several different methods, outlined in Appendix A. However, M. Shahbandeh et al. (2023) report PSF photometry for SN 2017eaw, and we find that our aperture photometry methods result in flux values for the Cycle 1 observations which are 10-40% higher than those reported in M. Shahbandeh et al. (2023). For consistency with the published photometry, we instead report PSF photometry done using space_phot333space_phot version 0.2.5 https://space-phot.readthedocs.io (J. Pierel, 2024; J. D. R. Pierel et al., 2024). We note that this choice does not significantly impact the conclusions of this work since they are based primarily on the difference between the JWST/MIRI epochs and not the absolute flux calibration. PSF photometry with space_phot is done on the stage 2 products for all filters except F2550W. This involves fitting the SN’s PSF in each of the 4 individual Level 2 CAL files using WebbPSF (M. D. Perrin et al., 2012, 2014, version 1.2.1) models. Given the low signal to noise detection of SN2017eaw in F2550W, we opt to do PSF photometry on the Level 3 stacked images for this filter. The space_phot routine for Level 3 photometry uses temporally and spatially dependent Level 2 PSF models from WebbPSF, and drizzles them together to create a Level 3 PSF model. While the MIRI PSF models have been significantly updated since the publication of M. Shahbandeh et al. (2023), we find our Cycle 1 PSF photometry is mostly consistent with the previously reported values (see Appendix A for more information). We report the flux values from space_phot for both the Cycle 1 and Cycle 2 observations in Table 1.
@@
| Filter | MJD | Phase444From explosion on 57885.2 MJD (S. D. Van Dyk et al., 2019) | Flux | AB Mag | MJD | Phase | Flux | AB Mag |
| [days] | [ mJy] | [days] | [ mJy] | |||||
@@ 2.3 MMT Optical and NIR Imaging
For Binospec observations, we utilize a standard dither pattern. These images are then reduced using a custom python Binospec imaging reduction pipeline555Initially written by K. Paterson and available on GitHub: https://github.com/CIERA-Transients/POTPyRI, which does standard flat-fielding, sky background estimation, astrometric alignments, and stacking of the final individual exposures.
For MMIRS, each observation consisted of a dithered sequence which alternates between the target field and a off-galaxy field to allow for better sky subtraction given the IR-brightness of NGC6946. The resulting , , and band observations were reduced using a custom pipeline666Adapted from the MMIRS imaging pipeline available on github: https://github.com/CIERA-Transients/POTPyRI which does standard dark-current correction, flat-fielding, sky background estimation and subtraction, astrometric alignments, and stacking of the final individual exposures.
The total field-of-view (FOV) of MMIRS (6’.96’.9) and Binospec (two 8’15’ FOVs with 3’ gap) is large enough to calibrate photometric zeropoints using isolated stars with cataloged Two Micron All Sky Survey (2MASS; M. F. Skrutskie et al., 2006) and Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; K. C. Chambers et al., 2016), respectively. We derive an effective (e)PSF model for each image by fitting bright, isolated stars with the EPSFBuilder tool from the photutils package in Astropy. For all filters where the supernova is detected, we then perform PSF-fitting at the location of the target as well as a set of 20 or more stars spread throughout the image. A low-order, two-dimensional polynomial is included in the fit to account for any spatially varying background and avoid over-fitting of the stars. To estimate the statistical uncertainty of each flux measurement, we first set the statistical uncertainty per pixel using the RMS error of the fit residuals scaled by a factor of the square root of the reduced (usually 1), then multiply by the number of ‘noise pixels’ of the ePSF777A derivation of this quantity by F. Masci can be found here: http://web.ipac.caltech.edu/staff/fmasci/home/mystats/noisepix_specs.pdf. We use the set of 2MASS or Pan-STARRS calibration stars to derive aperture corrections (0.1 mag in all filters) to scale PSF-fitting magnitudes to the images’ photometric zeropoints. We adopt the statistical flux uncertainty summed in quadrature with the RMS error of the stars used in the zeropoint and ePSF aperture correction as the total uncertainty in our reported magnitudes. Despite the large FOV of both MMIRS and Binospec, the limited number of isolated 2MASS and Pan-STARRS stars means that the zeropoint RMS dominates the reported error.
@@ 3.1 Comparison with Stars in the Field
We first identify bright IR objects with minimal variability in the field by using the 2MASS catalog (M. F. Skrutskie et al., 2006). These objects are then confirmed to be point sources in all of the JWST filters. Since star clusters may appear as point sources in MIRI filters, we also include a cut to remove any objects in crowded regions by using an HST F814W image of NGC6946 as a reference. This procedure results in 9 comparison stars for the SN 2017eaw field. Given the low signal to noise detections of these reference stars, particularly in the redder bands, we opt to do photometry on the stage 3 products for all filters rather than the stage 2 products as was done for SN 2017eaw. We note that stage 2 and 3 photometry of SN 2017eaw produce flux change measurements that are consistent within the uncertainties.
@@ 3.4 Dust Modeling Methods
Given the prominent shape of the silicate feature at m, the observed dust is unlikely to be optically thick. Therefore we first fit the dust using an optically thin model as was done in M. Shahbandeh et al. (2023) and S. Zsíros et al. (2024). We also model a dusty sphere (similar to M. Shahbandeh et al., 2023) and a dusty shell motivated by work presented in E. Dwek & R. G. Arendt (2024); for both of these dust models we allow the optical depth to vary. All of the models presented here assume there are two temperature components, as motivated by the SED shape, within the same geometry. Details of the luminosity equations for these three models can be found in Appendix B.
@@
As shown in Figure 2, the 5-25 m SED exhibits two distinct peaks with a trough at 13 m. This double humped shape is characteristic of optically thin silicate dust. Therefore we assume that the cool component is silicate dust, with g cm-3 and m as given by A. Laor & B. T. Draine (1993). The composition of the hot component is not as clearly identified. Recent studies report both amorphous carbon and silicate hot dust components around CCSNe (G. Hosseinzadeh et al., 2023b; M. Shahbandeh et al., 2023; S. Zsíros et al., 2024), so we try models with both silicate and amorphous carbon dust (m and from L. Colangeli et al., 1995) for the optically thin dust case. For the dusty sphere and dusty shell cases, we assume both components are silicate dust.
We attempt to fit a model of hot amorphous carbon dust and cool silicate optically thin dust, as was done for SN 1980K (S. Zsíros et al., 2024), to the NIR and JWST/MIRI photometry from Cycle 2. As shown in Figure 5, we find that the carbon+silicate model can not reproduce the photometry 15 m. We note that this excess could potentially be fit with an additional cooler component, as might be expected of dust formed in the ejecta or pre-existing dust at larger radii. However, given that the SED evolution redward of 15µm is not well constrained, we avoid implementing a third dust component. The hot carbon component requires a temperature of T K to reproduce the observed NIR peak, this is significantly above the condensation temperature of amorphous carbon dust (T K; K. Lodders & B. Fegley, 1997). We therefore deem it unlikely that the hot dust component in SN 2017eaw is primarily carbonaceous dust. However, we cannot rule out the possibility that there is both carbon and silicate dust in the ejecta, though the shape and temperature of the SED would suggest a high ratio of silicate to carbon dust if carbon dust is present.
Given the poorer quality fit of the carbon+silicate optically thin model, we focus primarily on models where both dust components are silicate. In Table 3, we list the model parameters, their priors, and their best-fit values (median and 1 equal-tailed credibility interval), for the optically thin, dusty sphere, and dusty shell models of the full considered SED for Cycle 1 and Cycle 2. The best-fit models for all three iterations, and the separate dust components, are shown in Figure 6. Since the Cycle 1 SED does not include F770W or NIR observations, we also fit the Cycle 2 SED excluding these filters. For completeness, we also present the results of the Cycle 2 fits excluding the NIR and F770W observations in Table 3 and discuss them further in Appendix C.

### v2

@@ 1 Introduction
Observations of high redshift galaxies have revealed significant amounts of dust in the early universe (D. P. Marrone et al., 2018; T. Hashimoto et al., 2019; J. Witstok et al., 2023; V. Markov et al., 2024; A. Nanni et al., 2025). The majority of this dust is likely associated with core-collapse supernovae (CCSNe; H. L. Morgan & M. G. Edmunds, 2003; R. Maiolino et al., 2004; C. Gall et al., 2011; R. Schneider & R. Maiolino, 2024). Models of high redshift supernovae (SNe) and star formation rates indicate that SNe would need to produce between M⊙ of dust per SN (P. Todini & A. Ferrara, 2001; A. Sarangi et al., 2018; R. Schneider & R. Maiolino, 2024) and observations of nearby SN remnants have revealed dust masses within this range (L. Dunne et al., 2003; H. L. Morgan et al., 2003; J. Rho et al., 2008; I. De Looze et al., 2017; J. Chastenet et al., 2022; F. D. Priestley et al., 2022). However, the vast majority of near- and mid-infrared studies of nearby SNe undertaken prior to JWST have revealed significantly lower dust masses (10-2 ; C. Gall et al., 2011; T. Szalai & J. Vinkó, 2013; T. Szalai et al., 2019a; S. Tinyanont et al., 2016). These previous studies therefore suggest that infrared observations of CCSNe in the decades after explosion may be missing a significant portion of the dust.
Dust formation in SNe likely occurs in the expanding ejecta interior to the reverse shock and/or in a cold dense shell between the forward and reverse shock created by the interaction between the forward shock and surrounding dense circumstellar material (CSM) (e.g. M. Pozzo et al., 2004; S. Mattila et al., 2008; N. Smith et al., 2008, 2009). In the decades to centuries following the SN, some of the dust in the interior ejecta will be destroyed by interaction with the reverse shock; however, some percentage is expected to survive this interaction and facilitate further dust formation (S. Bianchi & R. Schneider, 2007; T. Nozawa et al., 2007; D. W. Silvia et al., 2010). There are some indications that the majority of newly formed dust in the first years post-explosion in non-interacting hydrogen-rich CCSNe (e.g. type IIP/L SNe) is interior and optically thick and thus only visible in the spectral line profiles and not in infrared images. Studies modeling the optical nebular spectra of non-interacting H-rich SNe recover higher dust masses than indicated by infrared photometry alone (M. Niculescu-Duvaz et al., 2022; S. Zs´ıros et al., 2024). Further, the significant dust mass in SN 1987A is only observable in the far-IR and sub-millimeter (P. Bouchet et al., 2006; M. Matsuura et al., 2011; R. Indebetouw et al., 2014; M. Matsuura et al., 2015; P. Cigan et al., 2019). Thus even in the case where the dust is optically thin, any newly-formed dust may be too cold to detect in the near/mid-infrared.
@@
The launch of JWST has ushered in a new era of SN dust studies. JWST’s sensitivity, resolution and wavelength coverage has allowed for observations that probe both hot and cool dust around the SNe in the years following collapse. In concert with multi-wavelength ground- and space-based observations, numerous studies have utilized JWST data to constrain the dust formation around some of the nearest SNe (R. G. Arendt et al., 2023; G. Hosseinzadeh et al., 2023b; O. C. Jones et al., 2023; M. Shahbandeh et al., 2023; P. Bouchet et al., 2024; S. Gomez et al., 2024; M. Matsuura et al., 2024; M. Shahbandeh et al., 2024a, b; S. Zs´ıros et al., 2024; G. C. Clayton et al., 2025; A. Sarangi et al., 2025; T. Szalai et al., 2025; S. Tinyanont et al., 2025). This growing dataset is vital to understanding the onset of dust production in CCSNe and its impact on the evolution of the early Universe.
@@ 2 Observations
SN 2017eaw was discovered on 2017 May 14 in NGC 6946, a nearby galaxy ( Mpc according to the latest TRBG distance111https://edd.ifa.hawaii.edu/get_cmd.php?pgc=65001, this distance is also used in M. Shahbandeh et al. (2023)) with a high SN rate (see Figure 1). SN 2017eaw has extensive multi-wavelength observations, both pre- and post-explosion (explosion on 57885.2 MJD S. D. Van Dyk et al., 2019), and has been the subject of numerous studies (e.g. D. Y. Tsvetkov et al., 2018; J. Rho et al., 2018; C. D. Kilpatrick & R. J. Foley, 2018; S. Tinyanont et al., 2019; S. D. Van Dyk et al., 2019; T. Szalai et al., 2019b; L. Rui et al., 2019; R. J. Buta & W. C. Keel, 2019; K. E. Weil et al., 2020; M. Shahbandeh et al., 2023; K. A. Bostroem et al., 2023).
There is significant evidence of ejecta-CSM interaction in SN 2017eaw. Pre-explosion Spitzer Space Telescope (hereafter Spitzer) images show the progenitor star was surrounded by a dusty shell at 4000 (C. D. Kilpatrick & R. J. Foley, 2018). Early time detections of SN 2017eaw in the X-ray, UV, and radio indicate moderate interaction between the SN shock and the CSM (T. Szalai et al., 2019b). Further, the optical light curve of SN 2017eaw exhibits a bump peaking around a week post-explosion that further suggests early time CSM interaction (T. Szalai et al., 2019b). Years after explosion there remain signs of ongoing ejecta-CSM interaction. Hubble Space Telescope (HST) near-UV imaging reveals the SN is still UV bright (as shown in Figure 1; S. D. Van Dyk et al., 2023), light curves of the SN out to 1500 days indicate a shallower decline than expected for pure radioactive decay (M. Rizzo Smith et al., 2023), and late-time ( days post-explosion) optical spectra exhibit boxy line profiles indicating that the ejecta is continuing to collide with the surrounding material (K. E. Weil et al., 2020; M. Shahbandeh et al., 2023). Given the continued detection of CSM interaction, it is likely that some pre-existing dust surrounds SN 2017eaw.
@@ 2.1 JWST/MIRI
We attempt aperture photometry on the Cycle 1 and 2 images using several different methods, outlined in Appendix A. However, M. Shahbandeh et al. (2023) report PSF photometry for SN 2017eaw, and we find that our aperture photometry methods result in flux values for the Cycle 1 observations which are 10-40% higher than those reported in M. Shahbandeh et al. (2023). For consistency with the published photometry, we instead report PSF photometry done using space_phot222space_phot version 0.2.5 https://space-phot.readthedocs.io (J. Pierel, 2024; J. D. R. Pierel et al., 2024). We note that this choice does not significantly impact the conclusions of this work since they are based primarily on the difference between the JWST/MIRI epochs and not the absolute flux calibration. PSF photometry with space_phot is done on the stage 2 products for all filters except F2550W. This involves fitting the SN’s PSF in each of the 4 individual Level 2 CAL files using WebbPSF (M. D. Perrin et al., 2012, 2014, version 1.2.1) models. Given the low signal to noise detection of SN2017eaw in F2550W, we opt to do PSF photometry on the Level 3 stacked images for this filter. The space_phot routine for Level 3 photometry uses temporally and spatially dependent Level 2 PSF models from WebbPSF, and drizzles them together to create a Level 3 PSF model. While the MIRI PSF models have been significantly updated since the publication of M. Shahbandeh et al. (2023), we find our Cycle 1 PSF photometry is mostly consistent with the previously reported values (see Appendix A for more information). We report the flux values from space_phot for both the Cycle 1 and Cycle 2 observations in Table 1.
@@
| Filter | MJD | Phase333From explosion on 57885.2 MJD (S. D. Van Dyk et al., 2019) | Flux Density | AB Mag | MJD | Phase | Flux Density | AB Mag |
| [days] | [Jy] | [days] | [Jy] | |||||
@@ 2.3 MMT Optical and NIR Imaging
For Binospec observations, we utilize a standard dither pattern. These images are then reduced using a custom python Binospec imaging reduction pipeline444Initially written by K. Paterson and available on GitHub: https://github.com/CIERA-Transients/POTPyRI, which does standard flat-fielding, sky background estimation, astrometric alignments, and stacking of the final individual exposures.
For MMIRS, each observation consisted of a dithered sequence which alternates between the target field and a off-galaxy field to allow for better sky subtraction given the IR-brightness of NGC6946. The resulting , , and band observations were reduced using a custom pipeline555Adapted from the MMIRS imaging pipeline available on github: https://github.com/CIERA-Transients/POTPyRI which does standard dark-current correction, flat-fielding, sky background estimation and subtraction, astrometric alignments, and stacking of the final individual exposures.
The total field-of-view (FOV) of MMIRS (6’.96’.9) and Binospec (two 8’15’ FOVs with 3’ gap) is large enough to calibrate photometric zeropoints using isolated stars with cataloged Two Micron All Sky Survey (2MASS; M. F. Skrutskie et al., 2006) and Panoramic Survey Telescope and Rapid Response System (Pan-STARRS; K. C. Chambers et al., 2016), respectively. We derive an effective (e)PSF model for each image by fitting bright, isolated stars with the EPSFBuilder tool from the photutils package in Astropy. For all filters where the supernova is detected, we then perform PSF-fitting at the location of the target as well as a set of 20 or more stars spread throughout the image. A low-order, two-dimensional polynomial is included in the fit to account for any spatially varying background and avoid over-fitting of the stars. To estimate the statistical uncertainty of each flux measurement, we first set the statistical uncertainty per pixel using the RMS error of the fit residuals scaled by a factor of the square root of the reduced (usually 1), then multiply by the number of ‘noise pixels’ of the ePSF666A derivation of this quantity by F. Masci can be found here: http://web.ipac.caltech.edu/staff/fmasci/home/mystats/noisepix_specs.pdf. We use the set of 2MASS or Pan-STARRS calibration stars to derive aperture corrections (0.1 mag in all filters) to scale PSF-fitting magnitudes to the images’ photometric zeropoints. We adopt the statistical flux uncertainty summed in quadrature with the RMS error of the stars used in the zeropoint and ePSF aperture correction as the total uncertainty in our reported magnitudes. Despite the large FOV of both MMIRS and Binospec, the limited number of isolated 2MASS and Pan-STARRS stars means that the zeropoint RMS dominates the reported error.
@@ 3.1 Comparison with Stars in the Field
We first identify infrared bright objects with minimal variability in the field by using the 2MASS catalog (M. F. Skrutskie et al., 2006). These objects are then confirmed to be point sources in all of the JWST filters. Since star clusters may appear as point sources in MIRI filters, we also include a cut to remove any objects in crowded regions by using an HST F814W image of NGC6946 as a reference. This procedure results in 9 comparison stars for the SN 2017eaw field. Given the low signal to noise detections of these reference stars, particularly in the redder bands, we opt to do photometry on the stage 3 products for all filters rather than the stage 2 products as was done for SN 2017eaw. We note that stage 2 and 3 photometry of SN 2017eaw produce flux change measurements that are consistent within the uncertainties.
@@ 3.4 Dust Modeling Methods
Given the prominent shape of the silicate feature at m, the observed dust is unlikely to be optically thick. Therefore we first fit the dust using an optically thin model as was done in M. Shahbandeh et al. (2023) and S. Zs´ıros et al. (2024). We also model a dusty sphere (similar to M. Shahbandeh et al., 2023) and a dusty shell motivated by work presented in E. Dwek & R. G. Arendt (2024); for both of these dust models we allow the optical depth to vary. All of the models presented here assume there are two temperature components, as motivated by the SED shape, within the same geometry. Details of the luminosity equations for these three models can be found in Appendix B.
@@
As shown in Figure 2, the 5-25 m SED exhibits two distinct peaks with a trough at 13 m. This double humped shape is characteristic of optically thin silicate dust. Therefore we assume that the cool component is silicate dust, with g cm-3 and m as given by A. Laor & B. T. Draine (1993). The composition of the hot component is not as clearly identified. Recent studies report both amorphous carbon and silicate hot dust components around CCSNe (G. Hosseinzadeh et al., 2023b; M. Shahbandeh et al., 2023; S. Zs´ıros et al., 2024), so we try models with both silicate and amorphous carbon dust (m and from B. T. Draine 1985 and L. Colangeli et al. 1995 for silicate and amorphous carbon respectively) for the optically thin dust case. For the dusty sphere and dusty shell cases, we assume both components are silicate dust.
We attempt to fit a model of hot amorphous carbon dust and cool silicate optically thin dust, as was done for SN 1980K (S. Zs´ıros et al., 2024), to the NIR and JWST/MIRI photometry from Cycle 2. As shown in Figure 5, we find that the carbon+silicate model can not reproduce the photometry 15 m. We note that this excess could potentially be fit with an additional cooler component, as might be expected of dust formed in the ejecta or pre-existing dust at larger radii. However, given that the SED evolution redward of 15µm is not well constrained, we avoid implementing a third dust component. The hot carbon component requires a temperature of T K to reproduce the observed NIR peak, this is significantly above the condensation temperature of amorphous carbon dust (T K; K. Lodders & B. Fegley, 1997). We therefore deem it unlikely that the hot dust component in SN 2017eaw is primarily carbonaceous dust. However, we cannot rule out the possibility that there is both carbon and silicate dust in the ejecta, though the shape and temperature of the SED would suggest a high ratio of silicate to carbon dust if carbon dust is present.
Given the poorer quality fit of the carbon+silicate optically thin model, we focus primarily on models where both dust components are silicate. In Table 3, we list the model parameters, their priors, and their best-fit values (median and 1 equal-tailed credibility interval), for the optically thin, dusty sphere, and dusty shell models of the full considered SED for Cycle 1 and Cycle 2. The best-fit models for all three iterations, and the separate dust components, are shown in Figure 6 (the corresponding corner plots are available in the online journal as a figure set accompanying Figure 7). Since the Cycle 1 SED does not include F770W or NIR observations, we also fit the Cycle 2 SED excluding these filters. For completeness, we also present the results of the Cycle 2 fits excluding the NIR and F770W observations in Table 3 and discuss them further in Appendix C.